# From Cash to Cashless: UPI’s Impact on Spending Behavior among Indian Users

### v2

@@ 6. Conclusion
In this paper, we conducted a user study to understand the influence of UPI on spending behavior among Indian users. We studied how UPI has enabled convenience and seamless financial transactions, overshadowing cash transactions. UPI has played a pivotal and remarkable role in India’s push for a cashless economy and is also benefiting India’s GDP. Our study culminates by aligning findings with research questions and user recommendations, providing actionable insights for enhancing UPI app features and facilitating effective expenditure management. This research has broader implications for policymakers, businesses, and financial institutions, highlighting the need to adapt and promote responsible UPI usage to maximize its transformative potential in the Indian economic landscape.
Appendix A SURVEY QUESTIONS
1. What is your age?
2. What is your profession?
3. What is your gender? (Options: Male, Female, Non-binary, Prefer not to say)
4. Have you used UPI as the payment method?
5. If yes, which year did you start using it?
6. How often do you use UPI for making payments? (Daily, Weekly, Monthly, Occasionally, Rarely)
7. Have you noticed any changes in your spending habits since you started using UPI?
8. On a scale of 1 to 5, rate the convenience you find with UPI for making payments.
9. How satisfied are you with your overall experience using UPI for payments? (1 = Not Satisfied at all, 5 = Extremely Satisfied)
10. On average, how much do you spend daily using UPI?
11. Have you found any change in your spending ever since using UPI?
12. Do you feel that using UPI affects your ability to stick to your budget?
13. Have you experienced any instances where you have exceeded your budget or spent more than intended due to UPI usage?
14. How often do you check your transaction history or keep track of your UPI spending? (options: rarely, occasionally, frequently, very frequently, regularly)
15. How do you manage/track your expenses? (Which app do you use, if you do, e.g., split-wise, etc)
16. Additional Comments: Do you want to share anything about your experience with UPI and its impact on your spending behavior?
@@ 17. Will you be available for online interviews if called? (Yes/No)
18. If yes, provide your email address :
19. If yes, provide your phone number :
Appendix B INTERVIEW QUESTIONS
Demographic Information:
a. What is your age?
@@ Appendix B INTERVIEW QUESTIONS
b. What is your gender? (Options: Male, Female, Non-binary, Prefer not to say)
c. What is your occupation?
Usage of UPI:
a. Have you used UPI as a payment method? (Yes/No)
b. If yes, when did you start using it?
@@
c. How often do you use UPI to make payments? (Daily, Weekly, Monthly, Occasionally, Rarely)
@@
d. Can you briefly describe your experience with using UPI to make transactions?
@@
e. Have you noticed any changes in your spending habits since you started using UPI? (options: ”Increased spending,” ”Decreased spending,” or ”No change”)
@@
f. If yes, could you elaborate on the changes you have observed in your spending habits after adopting UPI?
Perception of UPI Convenience:
@@
a. On a scale of 1 to 5, rate the convenience you find with UPI for making payments. (1 = Not Convenient at all, 5 = Extremely Convenient)
b. In what ways do you perceive UPI to be convenient/inconvenient for you?
@@
c. On a scale of 1 to 5, how would you rate the convenience of using UPI compared to traditional payment methods like cash? (1 = Not Convenient at all, 5 = Extremely Convenient)
@@
d. How satisfied are you with your overall experience of using UPI for payments? (1 = Not Satisfied at all, 5 = Extremely Satisfied)
Income Information:
@@
a. What is your approximate annual income range?
Spending Behavior:
a. On average, what is the transaction amount when you use UPI?
b. On average, how much do you spend daily using UPI?
c. On average, how much do you spend monthly using UPI?
c1. Where do you spend your money?
@@
c2. Does the spending amount include only UPI transactions or all spending in general?
@@
d. Have you found any change in your spending ever since using UPI?
@@
e. If yes, could you specify the reasons for the increased spending after adopting UPI?
@@
f. Do you feel that using UPI affects your ability to stick to your budget? (Options: Strongly Disagree, Disagree, Neutral, Agree, Strongly Agree)
@@
g. How often do you check your transaction history or keep track of your UPI spending? (Options: Rarely, Occasionally, Frequently, Very Requently, Regularly)
@@
h. Have you experienced any instances where you have exceeded your budget or spent more than intended due to UPI usage?
@@
i. In your opinion, does using UPI encourage you to spend more money overall? (Options: Strongly Disagree, Disagree, Neutral, Agree, Strongly Agree)
Factors Influencing Spending Behavior:
@@
a. Besides UPI usage, do other factors have influenced your spending behavior?
b. If yes, please specify those factors.
@@
c. Are you aware of any psychological factors influencing your spending behavior when using UPI? If yes, please elaborate.
Attitudes Towards Financial Management:
@@
a. How would you describe your level of financial discipline? (1 to 5)
@@
b. Do you feel that using UPI has made managing your finances easier or more difficult? (“Easier,” “More Difficult,” “No Change”)
@@
c. How do you rate the statement ”Do you feel that UPI provides you with a clear understanding of your spending patterns”? (Options: Strongly Disagree, Disagree, Neutral, Agree, Strongly Agree)
@@
d. How do you rate the statement ”UPI has simplified your financial management”? (Options: Strongly Disagree, Disagree, Neutral, Agree, Strongly Agree)
Additional Comments:
@@
a. Is there anything else you would like to share about your experience with UPI and its impact on your spending behavior?

### v3

@@ 13. Conclusion
This study provides valuable insights into how UPI influences individual spending behavior and highlights user-driven improvements for UPI applications. Our mixed-methods approach revealed that a majority of users report increased spending due to UPI’s intangible nature, while also expressing high satisfaction with its convenience.
We developed a high-fidelity prototype based on user feedback, incorporating features like an expense tracker and budgeting tools. Usability testing confirmed that these features, particularly those promoting financial awareness, were highly valued by users. Our research offers actionable recommendations for UPI app developers and other stakeholders to enhance user engagement and support responsible financial management.
Appendix A PHASE 1 : SURVEY QUESTIONS
1. What is your age? (Open-ended)
2. What is your profession? (Multiple Choice) (Options: Student, Working Professional, Business, Others)
3. What is your gender? (Multiple Choice) (Options: Male, Female, Non-binary, Prefer not to say)
4. Have you used UPI as the payment method? (Yes/No)
5. If yes, which year did you start using it? (Open-ended)
6. How often do you use UPI for making payments? (Multiple Choice) (Options: Daily, Weekly, Monthly, Occasionally, Rarely)
7. Have you noticed any changes in your spending habits since you started using UPI? (Multiple Choice) (Options: Increased Spending, Decreased Spending, No Change)
8. On a scale of 1 to 5, rate the convenience you find with UPI for making payments. (Likert Scale)
9. How satisfied are you with your overall experience using UPI for payments? (Likert Scale) (1 = Not Satisfied at all, 5 = Extremely Satisfied)
10. On average, how much do you spend daily using UPI? (Open-ended)
11. Have you found any change in your spending ever since using UPI? (Yes/No)
12. Do you feel that using UPI affects your ability to stick to your budget? (Likert Scale) (Options: Strongly Disagree, Disagree, Neutral, Agree, Strongly Agree)
13. Have you experienced any instances where you have exceeded your budget or spent more than intended due to UPI usage? (Yes/No)
14. How often do you check your transaction history or keep track of your UPI spending? (Multiple Choice) (Options: Rarely, Occasionally, Frequently, Very Frequently, Regularly)
15. How do you manage/track your expenses? (Open-ended)
16. Additional Comments: Do you want to share anything about your experience with UPI and its impact on your spending behavior? (Open-ended)
@@ 17. Will you be available for online interviews if called? (Yes/No)
18. If yes, provide your email address: (Open-ended)
19. If yes, provide your phone number: (Open-ended)
Appendix B PHASE 1 : INTERVIEW QUESTIONS
Demographic Information
1. What is your age? (Open-ended)
@@
2. What is your gender? (Multiple Choice) (Options: Male, Female, Non-binary, Prefer not to say)
3. What is your occupation? (Open-ended)
Usage of UPI
4. Have you used UPI as a payment method? (Yes/No)
5. If yes, when did you start using it? (Open-ended)
@@
6. How often do you use UPI to make payments? (Multiple Choice) (Options: Daily, Weekly, Monthly, Occasionally, Rarely)
@@
7. Can you briefly describe your experience with using UPI to make transactions? (Open-ended)
@@
8. Have you noticed any changes in your spending habits since you started using UPI? (Multiple Choice) (Options: Increased Spending, Decreased Spending, No Change)
@@
9. If yes, could you elaborate on the changes you have observed in your spending habits after adopting UPI? (Open-ended)
Perception of UPI Convenience
@@
10. On a scale of 1 to 5, rate the convenience you find with UPI for making payments. (Likert Scale) (1 = Not Convenient at all, 5 = Extremely Convenient)
11. In what ways do you perceive UPI to be convenient/inconvenient for you? (Open-ended)
@@
12. On a scale of 1 to 5, how would you rate the convenience of using UPI compared to traditional payment methods like cash? (Likert Scale) (1 = Not Convenient at all, 5 = Extremely Convenient)
@@
13. How satisfied are you with your overall experience of using UPI for payments? (Likert Scale) (1 = Not Satisfied at all, 5 = Extremely Satisfied)
Income Information
@@
14. What is your approximate annual income range? (Open-ended)
Spending Behavior
15. On average, what is the transaction amount when you use UPI? (Open-ended)
16. On average, how much do you spend daily using UPI? (Open-ended)
17. On average, how much do you spend monthly using UPI? (Open-ended)
18. Where do you spend your money? (Open-ended)
@@
19. Does the spending amount include only UPI transactions or all spending in general? (Open-ended)
@@
20. Have you found any change in your spending ever since using UPI? (Yes/No)
@@
21. If yes, could you specify the reasons for the increased spending after adopting UPI? (Open-ended)
@@
22. Do you feel that using UPI affects your ability to stick to your budget? (Likert Scale) (Options: Strongly Disagree, Disagree, Neutral, Agree, Strongly Agree)
@@
23. How often do you check your transaction history or keep track of your UPI spending? (Multiple Choice) (Options: Rarely, Occasionally, Frequently, Very Frequently, Regularly)
@@
24. Have you experienced any instances where you have exceeded your budget or spent more than intended due to UPI usage? (Yes/No)
@@
25. In your opinion, does using UPI encourage you to spend more money overall? (Likert Scale) (Options: Strongly Disagree, Disagree, Neutral, Agree, Strongly Agree)
Factors Influencing Spending Behavior
@@
26. Besides UPI usage, do other factors have influenced your spending behavior? (Yes/No)
27. If yes, please specify those factors. (Open-ended)
@@ 26. Besides UPI usage, do other factors have influenced your spending behavior? (Yes/No)
28. Are you aware of any psychological factors influencing your spending behavior when using UPI? If yes, please elaborate. (Open-ended)
Attitudes Towards Financial Management
@@
29. How would you describe your level of financial discipline? (Likert Scale) (1 to 5)
@@
30. Do you feel that using UPI has made managing your finances easier or more difficult? (Multiple Choice) (Options: Easier, More Difficult, No Change)
@@
31. How do you rate the statement ”Do you feel that UPI provides you with a clear understanding of your spending patterns”? (Likert Scale) (Options: Strongly Disagree, Disagree, Neutral, Agree, Strongly Agree)
@@
32. How do you rate the statement ”UPI has simplified your financial management”? (Likert Scale) (Options: Strongly Disagree, Disagree, Neutral, Agree, Strongly Agree)
Additional Comments
@@
33. Is there anything else you would like to share about your experience with UPI and its impact on your spending behavior? (Open-ended)
Appendix C PHASE 2 : QUESTIONNAIRE
1. Age: (Open-ended)
2. Profession (Options: Student, Working Professional, Business, Others): (MCQ)
3. Gender (Options: Male, Female, Others): (MCQ)
4. Have you noticed any changes in your spending habits since you started using UPI? (Options: Increased Spending, Decreased Spending, Prefer not to disclose): (MCQ)
5. Which payment app do you use most often? (Options: PhonePe, Google Pay, Paytm, BHIM UPI, Others): (MCQ)
Expense Tracker
6. Do you believe adding an expense tracker to financial apps like PhonePe, Paytm, etc., would be beneficial? (Options: Not useful at all, Not useful, Neutral, Useful, Extremely Useful): (Likert scale)
7. Do you find it helpful to view expenses categorized by month and type (e.g., food, travel) with a day-wise breakdown and budget assigned to each category? (Options: Yes, it is very helpful; Yes, it is helpful but I don’t think it should be implemented; No, it is not helpful; No, it is confusing and unnecessary; Not able to answer): (MCQ)
8. Please provide any suggestions you have for improving the expense tracker: (Open-ended)
Balance Display
9. Do you think it is good to see your balance before and after making a payment? (Options: Yes, it is very helpful to understand and spend less; No, it does not add value and could be distracting; Not able to answer): (MCQ)
10. Does seeing your balance before payment help you reconsider large payments? (Options: Yes; No; Sometimes; Not able to answer): (MCQ)
Financial Management
11. Do you find the one-liner quotes about financial management, budgeting techniques, and related financial facts valuable? (Options: Yes, they can help me spend less and manage my finances better; Yes, they are informative but may not significantly impact my spending habits; No, they could be distracting and unlikely to influence my spending behavior; Not able to answer): (MCQ)
Budget Limitations
12. Do you think setting budget limits for specific categories (like food, travel) will help you manage your spending? (Options: Yes, it will help me spend less and manage my finances better; Yes, it might help, but I’m not sure how effective it will be; No, it likely won’t make a significant difference in my spending habits; No, I don’t think it will help with managing spending or reducing guilt; Not able to answer): (MCQ)
13. How often do you feel guilt or regret about overspending? (Options: Never, Rarely, Sometimes, Often, Always): (MCQ)
14. Describe a scenario where setting a budget limit for a specific category (e.g., food, travel) would be particularly beneficial for you: (Open-ended)
Payment Type Selection
15. Do you find it useful to select and label the type of payment (e.g., food, travel) either before or after proceeding with a transaction? (Options: Extremely useful, it helps me categorize and manage expenses better; Very useful, but it could be streamlined; Moderately useful, though it could be improved; Slightly useful, but it feels like unnecessary jargon; Not useful at all, it seems like a time waste; Not able to answer): (MCQ)
16. How can the feature to assign and label payment types be made more intuitive or useful? (Options: Simplify the labeling process to reduce unnecessary steps, Automatically categorize payments based on transaction history, Allow users to quickly select from frequently used categories, Provide an option to skip labeling for quick transactions, Use icons or visuals to represent different payment categories, Implement a drag-and-drop feature for categorizing expenses, Enable voice commands to assign and label payments, Provide personalized suggestions based on previous spending behavior, Not able to answer): (Multiple select)
Saving Streak Feature
17. Would you be motivated to save more if there were a saving streak feature that rewards you based on how much you save in different categories and overall? (Options: Yes, the rewards would motivate me to save more; No, I don’t think rewards would make a difference in my saving habits; Maybe, it depends on the type and value of the reward; Not able to answer): (MCQ)
18. What additional gamification features would you like to see on the app? (Options: Boosted cashback rewards for specific types of transactions or merchants, Incentives for making consecutive payments or transactions, Competitions that reward users for timely bill payments, Challenges that encourage sticking to budgeting goals with rewards, Exclusive, personalized offers based on spending patterns, Interactive scenarios to improve financial knowledge with immediate rewards, Random bonuses or discounts awarded for regular transactions, Not able to answer): (Multiple select)
Delay Feature
19. Would you find it helpful to have a timer (e.g., 8-10 seconds) before completing a large payment, allowing you time to reconsider and confirm if you really want to proceed? (Options: Yes, it would help me make more deliberate decisions and avoid impulsive spending; Yes, it might be useful, but I’m not sure if I’d use it frequently; No, it seems unnecessary and could be an inconvenience; No, I don’t think it would have an impact on my spending decisions; Not able to answer): (MCQ)
20. What do you consider a large payment in terms of rupees? (Open-ended)
Follow up Questions
21. Will you be available for online interviews? (Yes/No)
22. If yes, provide your email address: (Open-ended)
23. If yes, provide your phone number: (Open-ended)